OpenStudy (anonymous):

a is the point (0,-2) b is the point (-4,2) c is the point (0,,3) find an equation of the line that passes through C and is parallel to AB

OpenStudy (bibby):

use point-slope form using the slope of AB and the coordinates in C the slope fo rmula is \(\large m=\frac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

do you mean -2-2 over 0-4 =-4/4

OpenStudy (bibby):

\(x_1,y_1)\) is (0,-2) \(x_2,y_2)\) is (-4,2)

OpenStudy (bibby):

if we do it in your order it'd be -2-2 ---------- 0--4 (and not 0-4)

OpenStudy (anonymous):

-4/4 =-1 or gradient of -x

OpenStudy (bibby):

-1 is the slope of the line we want (parallel to the line containing A and B) now we plug it into the point slope form \(y-y_1=m(x-x_1)\)

OpenStudy (anonymous):

so y-3=-Cx-0)

OpenStudy (bibby):

"c is the point (0,,3)" I assume you meant y-3=-1(x-0)=y-3=-x

OpenStudy (anonymous):

yes

OpenStudy (bibby):

well get it into slope intercept form to finish

OpenStudy (anonymous):

how

OpenStudy (bibby):

y=mx+b

OpenStudy (bibby):

solve for y

OpenStudy (anonymous):

?

OpenStudy (anonymous):

whats the final answer

OpenStudy (bibby):

y-3=-x solve for y...

OpenStudy (anonymous):

y=3x???????

OpenStudy (bibby):

3-x or -x+3

OpenStudy (anonymous):

so the final answer is -x+3

OpenStudy (bibby):

yeah

OpenStudy (anonymous):

thank you very much